OpenStudy (anonymous):

If the sixth term of a sequence is 128 and the common ratio is 2, then what is the first term? 1 2 4 i think 2 but not shure

OpenStudy (anonymous):

an = a(r)^(n-1)

OpenStudy (anonymous):

um wat ?

OpenStudy (anonymous):

.hi im ur fann !!!

OpenStudy (campbell_st):

well you know n = 6, an = 128 and r = 2 so you have \[128 = a \times (2)^{6 -1}\] or \[128 = a \times 2^5\] just solve for a

OpenStudy (campbell_st):

which is the 1st term

OpenStudy (anonymous):

so a=4 ????????????

OpenStudy (campbell_st):

well that seems to work 2^5 = 32 and 32 x 4 = 128 well done